OpenStudy (narissa):

what is the solution to the equation? 2x+y=5

OpenStudy (narissa):

@Yuna123

OpenStudy (anonymous):

After doing some research and having some help from my family, here her is how I understand the equation and it's answer. You begin with 2x+y=5 ; then you try to isolate the variable being y. You subtract 2x, or honestly move it towards the other side to isolate y, and it goes to the other side ; making y=2x+5. I double checked it multiple times in different places due to most likely being near your age and learning as well, yet I believe this is the accurate answer. 2x+y=5 -2x -2x y=2x+5